OpenStudy (anonymous):

I need to convert an application problem into a linear equation. There is a pipe that is 240 in. cut into two pieces. One piece is three times the length of the other. Find the length of the pieces.

OpenStudy (anonymous):

3x + x = 240; x = 60in

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I need to find the length of the 2 pieces for this problem?

OpenStudy (anonymous):

is like you have x + y = 240. But y = 3x. Plug in and find both x and y. so, y = 60 * 3 = 180 in.

OpenStudy (anonymous):

I tried that, but when I put that answer in, the computer system said it was wrong.

OpenStudy (anonymous):

the small piece of pipe is 60 and the larger piece is 180

OpenStudy (anonymous):

I'll try that again, Thank You Very Much!!

OpenStudy (anonymous):

you are welcome.